Skip to Search

diagonal spacer

Systems Staff Meeting Notes
September 21, 2007

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Mark Sullivan
  1. Next weekend, Amy can start shooting stuff. She passed the Hunter's Ed course. The hunting season starts tomorrow.
  2. Discussion of phones - speed dialing, masking, InformaCast and parked calls.
  3. CNS Alerts - need to try again to drop Geoffrey's email and add Logan's.
Review Minutes of 9/7/07
  1. SW #6 - Lori asked for information on Patron Update changes.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders are flowing.
  1. Inventory - still having difficulties. LAC - we have not received anything at all. Special Collections - received in an unreadable format. Perhaps the inventory would be better if we required it be updated regularly. Perhaps a place to store it, like a shared directory? The goal is to keep them up to date. Maybe we can incorporate this into Grover? What aspect of it would we want to have opened to others?
  2. Verify Patron in West - Amy & Logan will be in this weekend to reconfigure DeepFreeze to allow updates. The problem is only in West.
  1. McAfee auto-updates are not working. Manual updates are working.
  2. Laptop batteries - Dell is going to send us 40 replacement batteries. They do not want the old ones in return. We will replace the batteries as they fail, not on mass.
  3. SAN - Logan expects to have the hardware configured by the end of next week. CNS is coming today to install the switch. The boxes in the vault need to be removed. Once it is up and running here, Logan will make time to go to MSL.
  4. While CNS is here today, we will talk to them about the phone in the server room.
  5. Public updates - we won't turn it on until the other public problem is resolved over the weekend.
  6. Bobbie is working with CNS on the wireless access in East. Public areas will be first.
Software Projects Status
  1. Patron Update request from Lori - just for information, no action taken.
  2. VIVO - spider web on linkages - ontology - Cornell thing. Not Institutional Repository.
  3. MSL - Val Davis & Sarah Russell-Gonzalez will request a Linux server as part of a mini-grant. It may be a virtual server.
  4. Problems with SQL 2K3 not deleting old backups.
  5. Grover - worked on security. Starting on ticket tab.
  6. Trudi - request for supervisors to see what training people under them have taken.
  7. Facilities - requested a change to LFPD to include PPD work order number.
  8. Evan has been added as an assignee to the Trouble Tickets.
  9. DLC Tracking - Erich has not tested v. 3.92. We will add one more feature to it and put it back up to test.
  10. UFDC - doing better - removed robots. The recovery is better. The backups are not failing as much. Yes, we want backup of the development server. We need to check on what exactly is getting backed up.
  11. Erich mentioned that he wanted more people (other than Mark) to know how to failover Greenstone. Mark to talk to Logan about it.
  12. Wiki and blogs - Logan and Stephen are working on it.
Web Projects
  1. Thanks to Logan and Will, the virtual Web server is up and running. Tom had to change the directory paths.
  2. Strategic Planning website is off the map.
  3. Fund raising web site is in the works.
  4. Other than that, Tom has been changing phone numbers and fixing pages.
Systems Security Management
  1. Meeting held.


University of Florida Home Page